Need help with setting up equation for...Mary bought 2 notebooks and 3 pens for $14.50. Dan bought 1 notebook and 2 pens for $7.50. How much does one notebook cost?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

Need help with setting up equation for...Mary bought 2 notebooks and 3 pens for $14.50. Dan bought 1 notebook and 2 pens for $7.50. How much does one notebook cost?
------
2n + 3p = 14.50
n + 2p = 7.50
--------------------
Modify for elimination::
2n + 3p = 14.50
2n + 4p = 1500
------------
Subtract and solve for "p"::
p = $0.50 (cost of one pen)
---
Solve for "n":
n + 2p = 7.50
n + 2*0.5 = 7.50
---
n + 1 = 7.50
n = $6.50 (cost of one notebook)
------
Cheers,
Stan H.
----------

Answer by Alan3354(69443)

What sort of help?
I can send you $5.